OpenStudy (anonymous):

What is the equation of the line passing through (0, 2) and perpendicular to the line y = 4x – 3 in slope-intercept form?

OpenStudy (anonymous):

perpendicular tells you slope is \[-\frac{1}{4}\]

OpenStudy (anonymous):

then you are done because you already have y intercept, so answer is \[y=-\frac{1}{4}x-2\]

OpenStudy (anonymous):

y = –4x + 3 y = 4x + 2 y = -1/4x + 2 y = 1/4x + 2

OpenStudy (anonymous):

Those Are the answer choices

OpenStudy (anonymous):

Satellite's answer is slightly incorrect. Since (0,2) is the intercept and -1/4 is the slope, and the formula for slope intercept form of a line is: y = mx + b where m is the slope and b is the y coordinate of the y intercept. Can you see his mistake?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

The answer is y=-1/4x+2. The y intercept (0,2) doesn't change. It is what it is. Only the slope changes for a perpendicular line.